OpenStudy (anonymous):

if you need to verify a trig function and are given cot^2x - cos^2x = cot^2xcos^2x how do you solver

OpenStudy (anonymous):

\[\cot^2 \theta -\cos^2 \theta = \cot^2 \theta \times \cos^2 \theta\]

OpenStudy (anonymous):

expand so it is all sines and cosines

OpenStudy (anonymous):

then work with one side and try to get the other side

OpenStudy (anonymous):

can you do it from here?

OpenStudy (anonymous):

\[\large RHS=\cot^2 \theta \times \cos^2 \theta\] \[\large =\frac{\cos^2 \theta}{\tan^2 \theta}\] \[\large =\cos^2 \theta \times \frac{\cos^2 \theta}{\sin^2 \theta}\] \[\large =(1-\sin^2 \theta)\times \frac{\cos^2 \theta}{\sin^2 \theta}\] \[\large =\frac{\cos^2 \theta}{\sin^2 \theta}-\cos^2 \theta\] \[\large =\cot^2 \theta -\cos^2 \theta\] \[\large =LHS\]

OpenStudy (anonymous):

Nice work @Azteck!